Two-point charges are attracted from each other with a force of 6kN. If the distance between them is reduced to one fourth of its original value, determine the new force of attraction between them.


1
Expert's answer
2021-10-12T12:33:00-0400

According to the Coulomb's law, the force between two point charges is proportional to the squared distance between them:


"F_1\sim\dfrac{1}{r_1^2}"

where "F_1 = 6kN" is the force and "r_1" is the distance in first case. Similarly:


"F_2\sim\dfrac{1}{(r_1/4)^2}"

where "F_2" is the new force and "r_1\/4" is the reduced distance. Taking the ration of these two expressions, obtain:


"\dfrac{F_2}{F_1} = \dfrac{r_1^2}{(r_1/4)^2}\\
F_2 = F_1\dfrac{r_1^2}{(r_1/4)^2} = \dfrac{F_1r_1^2\cdot 4^2}{r_1^2} = 16F_1\\
F_2 = 6kN\cdot 16 = 96kN"

Answer. 96kN.